OpenStudy (anonymous):

Factor completely p^4 – 18p^2 + 81

OpenStudy (phi):

let q = p^2 (so q^2 is (p^2)^2= p^4 ) you have q^2 -18q +81 can you factor this quadratic ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

what do you get ?

OpenStudy (anonymous):

(q-3)(q+3)

OpenStudy (phi):

?

OpenStudy (phi):

look at the 81: list pairs that multiply to give 81: 1,81 3,27 9,9 which pair will give 18 ? (the middle number , from -18q)

OpenStudy (anonymous):

my bad i meant to put 9

OpenStudy (phi):

ok, now what about the signs? the + sign on +81 means both roots are the same sign the - on the -18x means the largest root is negative

OpenStudy (anonymous):

so it would be (p-9)(p+9)

OpenStudy (phi):

no.... first it is still q (not p) second , the signs *are the same*

OpenStudy (anonymous):

okay, what would i do now?

OpenStudy (phi):

what do you have so far ?

OpenStudy (anonymous):

(q-9)(q-9)?

OpenStudy (phi):

yes... you can check by multiplying it out to see if you get the original... which you will. now we set q= p^2 . replace q with p^2. you get (p^2 - 9) ( p^2-9) can you factor (p^2-9) ?

OpenStudy (phi):

using "difference of squares" 9 is 3*3 so this is the same as (p^2 - 3^2)

OpenStudy (anonymous):

(p-3)(p+3), btw thanks man for teaching me not giving an answer

OpenStudy (phi):

yes, and we have two (p^2 - 3^2) : (p^2 - 3^2)(p^2 - 3^2) and you get (p-3)(p+3)(p-3)(p+3) you can write it a little more simply.... (p-3)(p-3)= (p-3)^2 and the same for (p+3)(p+3)

OpenStudy (phi):

either way is the answer, which one to use depends if you have multiple choice (p-3)(p+3)(p-3)(p+3) or (p-3)^2 (p+3)^2

OpenStudy (anonymous):

thanks